# Learning Unified Representation of 3D Gaussian Splatting

### v1

@@ Abstract
A well-designed vectorized representation is crucial for the learning systems natively based on 3D Gaussian Splatting. While 3DGS enables efficient and explicit 3D reconstruction, its parameter-based representation remains hard to learn as features, especially for neural-network-based models. Directly feeding raw Gaussian parameters into learning frameworks fails to address the non-unique and heterogeneous nature of the Gaussian parameterization, yielding highly data-dependent models. This challenge motivates us to explore a more principled approach to represent 3D Gaussian Splatting in neural networks that preserves the underlying color and geometric structure while enforcing unique mapping and channel homogeneity. In this paper, we propose an embedding representation of 3DGS based on continuous submanifold fields that encapsulate the intrinsic information of Gaussian primitives, thereby benefiting the learning of 3DGS.
@@ 1 Introduction
Despite its effectiveness in optimization-based reconstruction, we identify fundamental limitations when this parametric representation is employed as a learning space for neural networks. Our analysis reveals three critical issues. First, the mapping from parameters to rendered output exhibits non-uniqueness: multiple distinct parameter configurations yield identical visual results. This occurs through quaternion sign ambiguity (i.e., $\mathbf{q}$ and $\mathbf{q}$ represent identical rotations), symmetry-induced rotation invariance, and equivalences in spherical harmonic coefficients. Second, the parameter components exhibit severe numerical heterogeneity, with scales spanning large magnitude while quaternions remain unit-normalized. Third, these parameters inherently reside on distinct mathematical manifolds: positions in $\mathbb{R}^{3}$, rotations in $\text{SO}(3)$, and appearance in spherical harmonic space, inconsistent with the Euclidean space demands of standard neural architectures.
These theoretical limitations translate into substantial practical failures. Our empirical analysis reveals critical instabilities: negating quaternion values, which should produce equivalent rotations, causes complete decoding failure in parameter-trained autoencoders (see App. Fig. 8). Parametric representations achieve lower reconstruction quality and exhibit severe degradation under domain shift, particularly when transferring from object-level to scene-level data. Moreover, parametric embeddings lack robustness to noise, with minor perturbations causing disproportionate reconstruction errors. These consistent failures indicate fundamental representation deficiencies.
@@
We thus propose a principled alternative: representing each Gaussian primitive as a continuous field defined on its iso-probability surface. This submanifold field representation establishes a unique relationship between Gaussians and their geometric-photometric properties, eliminating the ambiguity inherent in parametric representations. By discretizing this field as a colored point cloud sampled from the probability surface, we obtain a representation that maintains uniform numerical properties and respects the underlying geometric structure. Particularly, we employ a variational autoencoder to learn embeddings from discretized submanifold fields, introducing a Manifold Distance metric based on optimal transport that better aligns with perceptual quality than parameter-space distances. Extensive experiments demonstrate substantial improvements in reconstruction quality, superior cross-domain generalization, and more robust latent representations compared to parametric baselines. The learned embeddings exhibit better noise resilience and semantic structure, validating our submanifold field representation as a more suitable learning space for 3D Gaussian primitives.
To summarize the contribution of this work, we:
@@ 2 Related Works
Embedding Gaussian Primitives. Recent works move beyond reconstruction and encode Gaussian parameters into latent spaces for tasks such as generation, editing, and compression.  Zhou et al. (2024b); Lin et al. (2025a); Wewer et al. (2024) learn structured latent variables from 3D Gaussian space to fulfill generation tasks. Editing and style-transfer methods use diffusion or style conditioning to manipulate Gaussians primitives in latent or rendering spaces (Chen et al., 2024a; Vachha & Haque, 2024; Lee et al., 2025; Palandra et al., 2024; Zhang et al., 2024; Kovács et al., 2024; Yu et al., 2024). Other works improve rendering quality by optimizing Gaussian parameters under diffusion priors (Tang et al., 2023; Yi et al., 2024; Chen et al., 2024d), while compression methods (Girish et al., 2024; Yang et al., 2025) reduce storage and computation by quantizing and embedding Gaussian parameters. These approaches show the potential of embedding Gaussians into neural latent spaces, but they assume Gaussian parameters are naturally compatible with neural learning, overlooking that these parameters were designed for optimization-based reconstruction. This oversight underlies our analysis in Section 3 and our proposal of a more suitable formulation.
@@ 3.2 Parameterization is Ill-Suited as a Learning Space
The non-uniqueness is from quaternion sign ambiguity, geometric symmetries, and rotation-spherical harmonic interactions producing equivalent parameter combinations (see Appendix for proof). The non-uniqueness of $\bm{\theta}$ will create “embedding collisions” where different parameter vectors produce identical rendered output (Wang & Isola, 2020). This makes the learning objective $\|\bm{\theta}_{\text{pred}}-\bm{\theta}_{\text{target}}\|_{p}$ ambiguous, as multiple parameter configurations can achieve the same visual result. The resulting conflicting gradients lead to training instability and poor convergence indicated by Bengio et al. (2013).
Numerical Heterogeneity The parameter components violate the homogeneous distribution assumption of standard neural architectures. Neural networks typically assume features share similar statistical properties for effective gradient flow (Ioffe & Szegedy, 2015). However, 3D Gaussian parameters span vastly different ranges. For example, pre-activation scales can range from 15-15 to $3$, while quaternions stay unit-normalized. More fundamentally, these parameters follow different distributions and live on different manifolds: positions $\bm{\mu}\in\mathbb{R}^{3}$, rotations $\mathbf{q}\in\text{SO}(3)$, scales $\mathbf{s}\in(\mathbb{R}^{+})^{3}$, and SH coefficients $c$ with exponential decay. Concatenating them ignores their heterogeneous nature. Small noises in quaternions can drastically alter geometry, while small noise in high-order SH coefficients is negligible, yet the network treats all dimensions equally.
@@ 4.1 Evaluation Setup
We implement and train three size-matched embedding models: our submanifold field VAE (Sec. 3.4), and two baseline parametric VAEs operating directly on $\bm{\theta}$. For the parametric models, each Gaussian primitive is represented as a 56-D vector (3+4+3K+13{+}4{+}3K{+}1 for $L_{\mathrm{max}}{=}3$), omitting global coordinates to match the SF-VAE setting. A three-layer MLP encodes this input to a 32-D latent ($56\!\!\rightarrow\!\!512\!\!\rightarrow\!\!512\!\!\rightarrow\!\!32\times 2$), and the decoder, either uses a MLP to map the latent back to $\widehat{\bm{\theta}}_{i}$, or uses the same decoder of SF-VAE to map to $\hat{\mathcal{P}}$. This setting further decouples evaluation results with the training objective functions. Apart from input dimension, all models share identical depth, width, latent size (32), and optimizer settings, ensuring a matched capacity.
@@
4.2 Quantitative Reconstruction Results
@@
We present a comprehensive quantitative and qualitative analysis of reconstruction quality for both object-level and scene-level data, as summarized in Tab. 4.2 and Fig. 3. All models are trained on the same randomly generated 3D Gaussian primitives dataset and evaluated on ShapeSplat and Mip-NeRF 360, using three matched encoder-decoder configurations to control for bias. Across all perceptual metrics (PSNR, SSIM, LPIPS), the submanifold field representation consistently outperforms parametric baselines. For example, on ShapeSplat, SF-VAE achieves substantially higher PSNR and SSIM and a much lower LPIPS, indicating both improved fidelity and perceptual quality. Similar performance gains are observed in scene-level reconstruction, where the submanifold field model demonstrates better performance across diverse spatial contexts.
Importantly, the Manifold Distance (M-Dist) metric shows a stronger empirical correlation with perceptual metrics like PSNR and LPIPS than traditional $L_{1}$ parameter distances, supporting our claim that M-Dist is a more robust and meaningful similarity measure for 3D Gaussian representations, truthfully reflecting perceptual differences rather than merely parameter discrepancies. The consistent improvement margin across both datasets highlights the advantage of learning in the submanifold field space, which better preserves intrinsic structure and view-dependent appearance, confirming the efficacy of our representation for high-fidelity 3D Gaussian modeling.
Figure 3: Qualitative results for rasterized reconstruction. Samples selected arbitrarily from Mip-NeRF 360 and ShapeSplat. Parametric models can induce confusion in rotation space, failing to predict the correct rotation.
4.3 Cross-Domain Generalization
Table 2: Results on representation generalization: train on one dataset, test on the other.
@@ 4.3 Cross-Domain Generalization
We evaluate cross-domain generalization by training on one dataset and testing on the other (object-level $\leftrightarrow$ scene-level) under an identical training protocol and capacity budget as in the reconstruction study. Concretely, we train either the proposed SF-VAE or the parametric MLP baselines on a source set and evaluate zero-shot on a target set, rendering novel views and reporting PSNR, SSIM, and LPIPS averaged over test samples (see Tab.  2).
Across both transfer directions, the submanifold field model consistently achieves higher reconstruction quality than the parametric baseline, with the largest margins when transferring from object-level sources to larger-scale scenes, indicating reduced sensitivity to dataset-specific statistics (scale, lighting, SH complexity).
4.4 Latent Space Evaluation
Robustness to Noise. To evaluate the submanifold field embedding’s robustness to noise, we gradually add higher levels of gaussian noise to the embedding space of each representation and test their reconstruction quality and M-Dist. As shown in Fig. 4, the embedding space of submanifold field model is more robust to random perturbation, this makes submanifold field embeddings a better learning target since it is less sensitive to potential systematic noise.
Figure 4: Parametric results use MLP encoder and SF-VAE’s decoder. Left: Sample visual comparison on robustness to noise in embedding space. Right: Comparison on M-Dist for different noise levels added to embedding space, tested on Mip-NeRF 360.
Latent Space Clustering. To further probe the semantic structure of the learned embedding spaces, we perform unsupervised graph clustering on both the raw Gaussian parameter space and the embedding outputs of each model. As visualized in Fig. 5, clusters formed in the submanifold field embedding space exhibit more detailed semantic separation against the reference images compared to those formed using normalized parameters or parametric embeddings. For example, SF-VAE’s embedding clustering in the first line of Fig. 5 outlines clearer separation of foreground objects with the background. The clusters appear smoother, less noisy, and with clearer boundaries, showing an ability to distinguish between different entities. This indicates that the submanifold field embedding captures more dense semantics and discriminative features, validating its usefulness.
Figure 5: Qualitative evaluation of unsupervised graph clustering, compared using Gaussian parameters and the embedding of the respective models, using the same hyperparameters. Submanifold field embeddings show better preservation of detailed semantics, validating its usefulness.
@@ 4.4 Latent Space Evaluation
4.5 More Studies
@@
Figure 6: Ablation study results, tested on Mip-NeRF 360. From left to right: (a) embedding space dimension, (b) generated training dataset size, (c) Submanifold Field sample size (all values normalized to $[0,1]$, with LPIPS inversely normalized).
@@
Latent Space Dimension. We evaluated different embedding space dimensions for the SF-VAE model to meet the best trade-off between compression and reconstruction quality. All models are trained on the generated dataset with $L=3$ order Spherical Harmonics. Results shown in Fig. 6 (a), 32 is the optimal balance point between reconstruction quality and latent space compression. All values are tested with baseline input/output of $P=12^{2}$. While this work does not specifically focus on compression effectiveness, embedding space robustness shown in Sec. 4.4 suggests potential in further latent tokenization and quantization.
@@
Size of Training Set. To empirically determine the number of random training samples required to achieve the best reconstruction results, we tested a wide spectrum of training sample sizes, ranging from 5K to 500K (baseline). The results indicate the proposed representation is data efficient. When only using 2% of the baseline training sample, our model can achieve close-to-baseline performance.
@@
Point Sample Size of Submanifold Fields. To ensure the submanifold fields are truthfully represented in a discrete manner, we evaluate different sample sizes $P$. Going from the lowest tested $P=6^{2}$ to the baseline $P=12^{2}$, we observe a gradual improvement in reconstruction quality, while going above $P=12^{2}$ yields very little improvement. Since $P$ directly correlates to the computational efficiency of the submanifold field model (see below), we keep $P=12^{2}$.
@@
Computational Efficiency. Increasing the point sample size $P$ leads to higher compute time and memory consumption. Encoding time stays low since a lightweight PointNet-based encoder shares weights with all input points, ensuring efficient encoding of large-scale scenes, with an inference speed of 1.72s per 1 million Gaussians for $P=12^{2}$ using a batch size of 4096 on an RTX 5090. Query time during decoding is $O(P)$ which is $O(n^{2})$ w.r.t. the grid size $n$. When the computation is not bottlenecked and uses a large enough batch size, the amortized time complexity is $\Theta(n)$.

### v2

@@ Abstract
A well-designed vectorized representation is crucial for the learning systems natively based on 3D Gaussian Splatting. While 3DGS enables efficient and explicit 3D reconstruction, its parameter-based representation remains hard to learn as features, especially for neural-network-based models. Directly feeding raw Gaussian parameters into learning frameworks fails to address the non-unique and heterogeneous nature of the Gaussian parameterization, yielding highly data-dependent models. This challenge motivates us to explore a more principled approach to represent 3D Gaussian Splatting in neural networks that preserves the underlying color and geometric structure while enforcing unique mapping and channel homogeneity. In this paper, we propose an embedding representation of 3DGS based on continuous submanifold fields that encapsulate the intrinsic information of Gaussian primitives, thereby benefiting the learning of 3DGS. Implementation available at https://github.com/cilix-ai/gs-embedding.
@@ 1 Introduction
Despite its effectiveness in optimization-based reconstruction, we identify fundamental limitations when this parametric representation is employed as a learning space for neural networks. Specifically, the native parameterization $\bm{\theta}$ conflicts with the inductive biases of standard neural architectures in three critical ways. First, the mapping from parameters to rendered output exhibits non-uniqueness. Ambiguities such as quaternion sign duality and symmetry-induced variances create a one-to-many mapping. This mapping creates “embedding collisions” where distinct parameter inputs producing identical visual outputs generate conflicting supervision signals, leading to training instability and poor convergence (Bengio et al., 2013; Wang and Isola, 2020). Second, the parameter components suffer from numerical heterogeneity. Spatial positions span large magnitudes while quaternions remain unit-normalized, violating the homogeneous feature distribution assumption required for effective gradient flow (Ioffe and Szegedy, 2015). Third, these parameters inherently reside on distinct mathematical manifolds, such as positions in $\mathbb{R}^{3}$, rotations in $SO(3)$, and appearance in spherical harmonic coefficients. Forcing these non-Euclidean variables into the Euclidean feature spaces of standard encoders breaks their intrinsic geometric structure, making the representation difficult to compress or regularize effectively.
These theoretical misalignments translate into substantial practical failures across various applications. Our empirical analysis reveals that parametric encoders are critically unstable; for instance, due to the sign ambiguity, simply negating a quaternion (a mathematically equivalent rotation) can cause complete decoding failure in parameter-trained autoencoders, see App. Fig. 9. This instability extends to generative tasks involving latent manipulation, where we observe that linear interpolation in the parametric latent space results in geometric “jitters” and unnatural transitions. Furthermore, parametric embeddings lack robustness to noise, with minor perturbations causing disproportionate reconstruction errors. Crucially, for downstream tasks extensively explored in recent work including generative modeling (Yi et al., 2024; Zhou et al., 2024b), compression (Shin et al., 2025; Girish et al., 2024), and editing (Chen et al., 2024a), these flaws manifest as discontinuous latent spaces and inefficient encoding. Instead of capturing robust geometric semantics, models are forced to resolve parameter ambiguities with more efforts, leading to sub-optimal performance in learning-based reconstruction (Charatan et al., 2024) and limited generalization across diverse domains.
@@
We thus propose a principled alternative that represents each Gaussian primitive as a continuous field defined on its iso-probability surface. This submanifold field representation establishes a unique correspondence between Gaussians and their geometric-photometric properties, removing the ambiguities of parametric representations. By discretizing this field as a colored point cloud sampled from the probability surface, we obtain a numerically uniform and geometrically consistent representation. We further employ a variational autoencoder to learn embeddings from these discretized submanifold fields, together with a Manifold Distance metric based on optimal transport that better correlates with perceptual quality than parameter-space distances. Extensive experiments show higher reconstruction quality, stronger cross-domain generalization, and more robust latent representations. The learned embedding space is smooth and generalizable, indicating strong potential for semantic understanding and generative modeling, as demonstrated through unsupervised segmentation and neural field decoding with the proposed embeddings. To summarize the contribution of this work, we:
@@ 2 Related Works
Embedding Gaussian Primitives. Recent works move beyond reconstruction and encode Gaussian parameters into latent spaces for tasks such as generation, editing, and compression.  Zhou et al. (2024b); Lin et al. (2025a); Wewer et al. (2024) learn structured latent variables from 3D Gaussian space to fulfill generation tasks. Editing and style-transfer methods use diffusion or style conditioning to manipulate Gaussians primitives in latent or rendering spaces (Chen et al., 2024a; Vachha and Haque, 2024; Lee et al., 2025; Palandra et al., 2024; Zhang et al., 2024; Kovács et al., 2024; Yu et al., 2024). Other works improve rendering quality by optimizing Gaussian parameters under diffusion priors (Tang et al., 2023; Yi et al., 2024; Chen et al., 2024d), while compression methods (Girish et al., 2024; Yang et al., 2025) reduce storage and computation by quantizing and embedding Gaussian parameters. These approaches show the potential of embedding Gaussians into neural latent spaces, but they assume Gaussian parameters are naturally compatible with neural learning, overlooking that these parameters were designed for optimization-based reconstruction. This oversight underlies our analysis in Section 3 and our proposal of a more suitable formulation.
@@ 3.2 Parameterization is Ill-Suited as a Learning Space
The non-uniqueness is from quaternion sign ambiguity, geometric symmetries, and rotation-spherical harmonic interactions producing equivalent parameter combinations (see App. A for proof). The non-uniqueness of $\bm{\theta}$ will create “embedding collisions” where different parameter vectors produce identical rendered output (Wang and Isola, 2020). This makes the learning objective $\|\bm{\theta}_{\text{pred}}-\bm{\theta}_{\text{target}}\|_{p}$ ambiguous, as multiple parameter configurations can achieve the same visual result. The resulting conflicting gradients lead to training instability and poor convergence indicated by Bengio et al. (2013).
Numerical Heterogeneity The parameter components violate the homogeneous distribution assumption of standard neural architectures. Neural networks typically assume features share similar statistical properties for effective gradient flow (Ioffe and Szegedy, 2015). However, 3D Gaussian parameters span vastly different ranges. For example, pre-activation scales can range from 15-15 to $3$, while quaternions stay unit-normalized. More fundamentally, these parameters follow different distributions and live on different manifolds: positions $\bm{\mu}\in\mathbb{R}^{3}$, rotations $\mathbf{q}\in\text{SO}(3)$, scales $\mathbf{s}\in(\mathbb{R}^{+})^{3}$, and SH coefficients $c$ with exponential decay. Concatenating them ignores their heterogeneous nature. Small noises in quaternions can drastically alter geometry, while small noise in high-order SH coefficients is negligible, yet the network treats all dimensions equally.
@@ 4.1 Evaluation Setup
We implement and train three size-matched embedding models: our submanifold field VAE (Sec. 3.4), and two baseline parametric VAEs operating directly on $\bm{\theta}$. For the parametric models, each Gaussian primitive is represented as a 56-D vector (3+4+3K+13{+}4{+}3K{+}1 for $L_{\mathrm{max}}{=}3$), omitting global coordinates to match the SF-VAE setting. A three-layer MLP encodes this input to a 32-D latent ($56\!\!\rightarrow\!\!512\!\!\rightarrow\!\!512\!\!\rightarrow\!\!32\times 2$), and the decoder, either uses a MLP to map the latent back to $\widehat{\bm{\theta}}_{i}$, or uses the same decoder of SF-VAE to map to $\hat{\mathcal{P}}$. This setting further decouples evaluation results with the training objective functions. Apart from input dimension, all models share identical depth, width, latent size (32), and optimizer settings (all using Adam), ensuring a matched capacity.
@@
4.2 Evaluation on Representation Learning Framework
@@
Zero-shot Reconstruction. We present a comprehensive quantitative and qualitative analysis of reconstruction quality for both object-level and scene-level data, as summarized in Tab. 1 and Fig. 3. All models are trained on the same randomly generated 3D Gaussian primitives dataset and evaluated on ShapeSplat and Mip-NeRF 360, using three matched encoder-decoder configurations to control for bias. Across all perceptual metrics (PSNR, SSIM, LPIPS), the submanifold field representation consistently outperforms parametric baselines. For example, on ShapeSplat, SF-VAE achieves substantially higher PSNR and SSIM and a much lower LPIPS, indicating both improved fidelity and perceptual quality. Similar performance gains are observed in scene-level reconstruction, where the submanifold field model demonstrates better performance across diverse spatial contexts.
Importantly, the Manifold Distance (M-Dist) metric shows a stronger empirical correlation with quality metrics like PSNR and LPIPS than traditional $L_{1}$ parameter distances, supporting our claim that M-Dist is a more robust and meaningful similarity measure for 3D Gaussian representations, truthfully reflecting perceptual differences rather than merely parameter discrepancies. The consistent improvement margin across both datasets highlights the advantage of learning in the submanifold field space, which better preserves intrinsic structure and view-dependent appearance, confirming the efficacy of our representation for high-fidelity 3D Gaussian modeling.
Figure 3: Qualitative results for rasterized reconstruction. Samples selected arbitrarily from Mip-NeRF 360 and ShapeSplat. Parametric models can induce confusion in parameter space, failing to embed and restore the correct Gaussian parameters.
Table 2: Reconstruction quality comparison under cross-domain setting. All models trained on either ShapeSplat or Mip-NeRF 360 dataset are tested on another dataset. We show that the generalization ability of SF Embedding framework is inherently domain-agnostic even without random data.
@@
Cross-domain Reconstruction. We also evaluate cross-domain generalization by training on one real-world dataset and testing on the other (object-level $\leftrightarrow$ scene-level) under an identical training protocol and capacity budget as in the reconstruction study. Concretely, we train either the proposed SF-VAE or the parametric MLP baselines on a source set and evaluate on a target set, rendering novel views and reporting PSNR, SSIM, and LPIPS averaged over test samples (see Tab.  2). Across both transfer directions, the SF-based embedding consistently achieves higher reconstruction quality than the parametric baseline, indicating reduced sensitivity to domain-specific statistics (e.g., scale, lighting, SH complexity). Particularly, comparing these transfer results with Tab. 1, we find that the model trained on synthetic random data actually outperforms the models transferred from real-world domains. This indicates that our random learning strategy effectively strips away domain priors, which establishes our approach as a unified representation, where a single model trained on synthetic data can fundamentally generalize to real-world contexts.
4.3 Sensitivity Study of Representation
Robustness to Noise. To evaluate the submanifold field embedding’s robustness to noise, we gradually add higher levels of gaussian noise to the embedding space of the parametric model and the submanifold field model and test their reconstruction quality and M-Dist. To ensure fair comparison, we use noise level as a ratio to variance instead of absolute noise magnitude. As shown in Fig. 4, the embedding space of submanifold field model is more robust to random perturbation, this makes submanifold field embeddings a better learning target since it is less sensitive to potential noise introduced by downstream regression.
Figure 4: Reconstruction results using embeddings with noise. Left: Visualization of reconstructed scene from noisy embeddings of Gaussian parameters (MLP) and SF-VAE. Right: Comparison on M-Dist for different noise levels added to embedding space, tested on Mip-NeRF 360. Noise level is defined as the ratio between the noise magnitude and the embedding variance.
@@ 4.3 Sensitivity Study of Representation
4.4 Representation Applicability
Unsupervised Clustering. To further probe the semantic structure of the learned embedding spaces, we perform unsupervised graph clustering on both the raw Gaussian parameter space and the embedding outputs of each model. As visualized in Fig. 5, clusters formed in the submanifold field embedding space exhibit more detailed semantic separation against the reference images compared to those formed using normalized parameters or parametric embeddings. For example, SF-VAE’s embedding clustering in the first line of Fig. 5 outlines clearer separation of foreground objects with the background. The clusters appear smoother, less noisy, and with clearer boundaries, showing an ability to distinguish between different entities. This indicates that the submanifold field embedding captures more dense semantics and discriminative features, validating its usefulness.
Figure 5: Unsupervised graph clustering based on raw Gaussian parameters and various embeddings. Submanifold field embeddings show better preservation of detailed semantics, showing its downstream applicability.
Gaussian Neural Fields. To validate the potential of our representation for advanced downstream tasks, we introduce the Gaussian Neural Field (GNF). Drawing inspiration from the decoding structures in generative diffusion models (e.g., DiffGS by Zhou et al. (2024b)) and neural compression frameworks (Wu and Tuytelaars, 2024), the GNF functions as a coordinate-based neural implicit field as illustrated in Fig. 6. Specifically, it employs a lightweight MLP (architecture detailed in App. D.4) to learn a continuous mapping from spatial coordinates $\mathbf{x}_{i}$ to per-primitive descriptors. This setup allows us to evaluate the “learnability” of our representation: while regressing heterogeneous raw parameters $\bm{\theta}_{i}$ often leads to optimization difficulties, our unified SF embeddings provide a smooth and well-conditioned target for the neural field. As evidenced in Tab. 3 and visualization in App. D.4, the SF-guided GNF outperforms the parameter-based baseline in visual fidelity with equivalent training effort. This indicates that our representation is more friendly to neural networks, hinting at its utility for potential downstream generative and compression tasks.
Figure 6: Setting of a Gaussian Neural Field, we compare between the prediction target SF embedding and raw GS parameters.
Table 3: Comparison between Gaussian Neural Fields trained using submanifold field embeddings and raw Gaussian parameters. Top: ShapeSplat, bottom: Mip-NeRF 360.
Target PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # Params
Raw GS Parameter 51.660 0.925 0.141 0.21M
SF Embedding 58.619 0.980 0.043 0.20M
Raw GS Parameter 19.922 0.648 0.410 1.87M
SF Embedding 24.395 0.804 0.261 1.85M
4.5 More Studies on Implementation Details
@@
Figure 7: Behavior studies tested on Mip-NeRF 360. From left to right: (a) embedding dimension, (b) generated training dataset size, (c) Submanifold Field discretized (i.e., point sample) grid size.
[…]
@@
Behavior Study of Latent Space Dimension. We evaluated different embedding space dimensions for the SF-VAE model to meet the best trade-off between compression and reconstruction quality. All models are trained on the generated dataset with $L=3$ order Spherical Harmonics. Results shown in Fig. 7 (a), 32 is the optimal balance point between reconstruction quality and latent space compression. All values are tested with baseline input/output of $P=12^{2}$. While this work does not specifically focus on compression effectiveness, embedding space robustness shown in Sec. 4.3 suggests potential in further latent tokenization and quantization.
@@
Behavior Study of Training Set Size. To determine the number of random training samples required to achieve the best reconstruction results, we vary the trainiing sample size from 5K to 500K (baseline), see Fig. 7 (b). The results indicate the proposed representation is data efficient. When only using 2% of the baseline training sample, our model can achieve close-to-baseline performance.
@@
Behavior Study of SF Discretized Size. To ensure the submanifold fields are truthfully represented in a discrete manner, we evaluate different sample sizes $P$, see Fig. 7 (c). Going from the lowest tested $P=6^{2}$ to the baseline $P=12^{2}$, we observe a gradual improvement in reconstruction quality, while going above $P=12^{2}$ yields very little improvement. Since $P$ directly correlates to the computational efficiency of the submanifold field model (see below), we keep $P=12^{2}$.
@@
Computational Efficiency. Increasing the point sample size $P$ increases computation and memory. Encoding time remains low since a lightweight PointNet-based encoder shares weights with all input points, giving an inference speed of 1.72s per 1 million Gaussians for $P=12^{2}$ with a batch size of 4096 on an RTX 5090. Decoding time is 4.20s per 1 million Gaussians, from embedding to Gaussian parameters. The complexity is $O(P)$ or $O(n^{2})$ w.r.t. the grid size $n$. We utilize the advantage of GPU parallel computation to boost the calculation for parameter fitting module. In the 4.20s of decoding time, the fitting module (PCA + SH fitting) consumes only about 0.48s which is negligible for large Gaussian scenes.